# Background Semantics Matter: Cross-Task Feature Exchange Network for Clustered Infrared Small Target Detection With Sky-Annotated Dataset

### v2

@@ V-A Experimental Settings
DETR [63] ResNet50 24.940G 41.555M 0.000 0.000 0.000 0.000
Sparse R-CNN [64] ResNet50 45.274G 0.106G 0.183 0.572 0.154 0.614
Deformable DETR [65] ResNet50 51.772G 40.099M 0.024 0.016 0.018 0.197
DAB-DETR [67] ResNet50 28.939G 43.702M 0.005 0.054 0.000 0.001
$\star$ BAFE-Net (Ours) ResNet18 57.654G 22.31M 0.283 0.335 0.233 0.325
Figure 6: Feature visualization comparison. Compared with other methods, the proposed BAFE-Net shows superior performance with higher detection of small objects and fewer false alarms.
V-B Comparison with State-of-the-Arts
We benchmarked our BAFE-Net against current leading or state-of-the-art methods on the proposed DenseSIRST dataset, using metrics such as $mAP_{07}$, $mAP_{12}$, and Recall. The experimental results are presented in Table II.
End-to-end (End2End) methods, while theoretically appealing due to their NMS-free style, demonstrated suboptimal performance. This can be attributed to their inherent requirement for extensive training data, akin to COCO-level volumes, for effective operation. The relatively limited availability of infrared small target detection datasets inherently handicaps such methods, resulting in their diminished performance.
In contrast, one-stage methods, which directly regress bounding boxes and categories on the feature pyramids, demonstrated superior performance. These methods are not only faster and more computationally efficient, but also achieved higher mAP and recall, indicating their overall effectiveness. Significantly, our BAFE-Net method outperformed all other one-stage methods. The superior performance of BAFE-Net validates the efficacy of our approach in accurately detecting small infrared targets amidst dense conditions.
TABLE III: Quantitative ablation study evaluates the impact of the Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting within the data augmentation module of BAFE-Net
Strategy Module DenseSIRST
Copy-Paste Cluster Gaussian Background-Aware mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) 0.225 0.311 0.180 0.314
(b) $\checkmark$ 0.236 0.322 0.203 0.303
(c) $\checkmark$ $\checkmark$ 0.242 0.329 0.205 0.335
(d) $\checkmark$ $\checkmark$ 0.245 0.321 0.211 0.338
(e) $\checkmark$ $\checkmark$ $\checkmark$ 0.250 0.319 0.215 0.315
(f) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.258 0.340 0.220 0.340
@@ V-B Comparison with State-of-the-Arts
TABLE IV: Quantitative ablation study evaluates the impact of the Background-Aware segmentation, Feature Exchange module, and adapter module within the Background-Aware Feature Exchange Network (BAFE-Net)
@@
TABLE V: Quantitative ablation study was conducted to evaluate the impact of the number of hidden layers in the Adapter
@@
  (a) 8 0.222 0.305 0.206 0.308
(b) 16 0.243 0.318 0.214 0.310
(c) 32 0.256 0.329 0.218 0.312
(d) 64 0.266 0.336 0.230 0.323
TABLE VI: Quantitative ablation study evaluates the impact of different Feature Exchange Mechanisms within the Feature Exchange module of BAFE-Net
@@
TABLE VII: Ablation study on the impact of feature selection on detection performance in Channel-wise Feature Exchange
@@
TABLE VIII: Ablation study on the impact of the number of exchange channels on detection performance in Top-K selection
@@
(h) 1/16 0.245 0.326 0.213 0.301
(i) 1/32 0.236 0.319 0.212 0.312
TABLE IX: Quantitative ablation study evaluates the impact of different backbones on BAFE-Net performance
Strategy ResNet DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) ResNet18 0.283 0.335 0.233 0.341
(b) ResNet34 0.281 0.354 0.226 0.332
(c) ResNet50 0.274 0.342 0.248 0.338
(d) ResNet101 0.261 0.317 0.241 0.336
@@ V-C Visual Analysis
Second, our method generates cleaner background activation maps compared to the other methods. As evident from Fig. 6, the background regions in our activation maps exhibit consistently lower activation intensities, effectively suppressing background interference. This can be attributed to the background-aware mechanism employed by BAFE-Net, which learns to differentiate between target and background regions during the feature exchange process. By minimizing the activation levels in the background areas, our method demonstrates enhanced robustness against false positives, leading to a higher detection precision.
V-D Ablation Study
V-D1 Ablation Study on Background-Aware Gaussian Copy-Paste
We initiate our investigation from the data augmentation module, focusing on the progressive integration of our proposed Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting. The experimental outcomes, as outlined in Table III, generate several key insights, providing a holistic understanding of the impact of each module. Firstly, the effective enhancement of detection performance by each module is noticeable, corroborating the initial motivation of our study. The Copy-Paste technique, even when used independently, significantly boosts performance, as evidenced by the increases of 1.1% and 2.3% on the DenseSIRST dataset, when comparing strategy (a) with strategy (b).
Further refinement through clustering of pasted objects also yields substantial improvements. This is evident when comparing strategies (b) to (c), and (d) to (e), respectively, indicating the considerable value of object clustering. Applying Gaussian blur to the edges of the pasted objects, as a means to enhance their realistic appearance, also proves to be significantly beneficial in improving detection performance. This is substantiated when comparing strategies (b) with (d), and (c) with (e).
Finally, the culmination of all modules in a Background-Aware Gaussian Copy-Paste context, as depicted by strategy (f), results in further enhancements in detection capability. This is demonstrated with an additional increase of 0.8% on the DenseSIRST dataset when comparing strategy (e) with strategy (f).
V-D2 Ablation Study on the Scheme of Background-Aware Feature Exchange
The Background-Aware Feature Exchange module constitutes the core contribution of this paper, as evidenced by its significant impact on detection performance demonstrated in Table IV. To validate the effectiveness of the proposed components, we conducted an extensive ablation study by exploring various combinations of the background-based segmentation, Feature Exchange module, and adapter module.
Table IV encapsulates the results of our experimental study and yields several crucial insights. Firstly, each constituent module substantively and independently enhances the detection performance, thereby validating the underpinnings of our research. Secondly, comparison between Strategy (a) and Strategy (b) reveals a significant performance augmentation when background-based segmentation is incorporated into object detection, wherein the detection head interacts with the segmentation head. This integration accounted for an increase of 1.0% and 0.3% on the DenseSIRST dataset.
Further, juxtaposition of Strategy (b) and Strategy (c) highlights the contribution of the Feature Exchange in the interaction mode between the detection head and the segmentation head. This interaction led to a notable improvement in detection performance by 1.1% on the DenseSIRST dataset. Upon comparing Strategy (c) and Strategy (d), it becomes evident that introducing feature adapation prior to the Feature Exchange further amplifies detection performance.
As the feature adapter employs a Multi-Layer Perceptron (MLP), the number of hidden layers in the MLP is a crucial parameter that warrants careful consideration. Increasing the number of hidden layers typically improves performance but comes at the cost of increased computational complexity. To strike a balance between performance and efficiency, we conducted an experiment to determine the optimal number of hidden layers, as shown in Table  V. The experimental results demonstrate that increasing the number of hidden layers leads to a consistent improvement in detection performance. However, the performance gain becomes less significant when the number of hidden layers exceeds 64, while the parameter count increases substantially. Therefore, we opted for 64 hidden layers in the Adapter, as it provides a significant performance boost while maintaining a reasonable computational overhead compared to using 128 hidden layers or more.
V-D3 Ablation Study on Feature Exchange Mechanism
In this part, a rigorous exploration of various Feature Exchange Mechanisms was undertaken. This encompassed diverse mechanisms such as Channel Exchange and Spatial Exchange, along with a multitude of their combinations. The quest was to ascertain the most efficacious method for Feature Exchange. Channel Exchange, as a mechanism, facilitates the direct interplay at the channel level amongst different features. In contrast, Spatial Exchange is implemented by orchestrating exchanges along the spatial dimensions inherent in the feature maps. The exchange ratio was judiciously set at 0.5 to maintain a balance.
As delineated in Table VI, the superiority of the Channel Exchange method over alternative Feature Exchange mechanisms is clearly manifested. The empirical evidence underscores the robustness and optimality of Channel Exchange, thereby validating its adoption as the preferred method for Feature Exchange in our model.
V-D4 Ablation Study in Dynamic Channel Selection
In this section, we first conduct a rigorous exploration of feature selection methods. This includes random selection, fixed selection, and dynamic selection. The task is to determine the most effective feature selection method. 1) Randomly Selection: This method selects feature channels randomly, which may lead to suboptimal feature exchanges. In some cases, it could even select features that hinder the task, ultimately reducing model performance. 2) Fixed Selection: While fixed selection provides consistency during training and eliminates uncertainty, it lacks the adaptability to dynamic input features, thus limiting the model’s performance in complex or varied scenarios. 3) Dynamic Selection: dynamic selection maximizes the performance of the model and reduces unnecessary information redundancy by selecting the most useful feature channels for exchange. The exchange ratio is judiciously set to 0.5 to maintain a balance.
As shown in Table VII, the advantages of dynamic selection are obvious. Empirical evidence highlights the robustness and optimality of dynamic selection, thus confirming its adoption as the preferred method for feature selection in our model.
In addition, we conduct a rigorous exploration of the value of $p$ in Top-K selection. As shown in Table VIII, $p$ is optimal when it is 1/2.
V-D5 Impact of Backbone Depth
For infrared small target detection, the depth of the backbone network emerges as a critical factor. A shallower network, while computationally efficient, might extract features with insufficient semantic richness, thereby impinging on the overall detection performance. Conversely, an excessively deep network might overwhelm the delicate features of smaller targets amidst the background, consequently leading to a performance degradation.
We conducted a series of experiments using different ResNet architectures, varying the depth from ResNet18 to ResNet101, to identify the optimal backbone for our proposed BAFE-Net. The results of these experiments are summarized in Table IX. Notably, ResNet18 has the highest $mAP_{07}$ metric and ResNet50 has the highest $mAP_{12}$ metric.

### v3

@@ IV-B2 Cross-Task Feature Hard-Exchange
DETR [24] ResNet50 24.940G 41.555M 0.000 0.000 0.000 0.000 0.000 0.001 0.000 0.001
Sparse R-CNN [25] ResNet50 45.274G 0.106G 0.136 0.370 0.160 0.388 0.826 0.953 0.827 0.952
Deformable DETR [26] ResNet50 51.772G 40.099M 0.091 0.018 0.015 0.126 0.551 0.782 0.550 0.770
Conditional DETR [27] ResNet50 27.143G 40.297M 0.000 0.001 0.000 0.002 0.000 0.010 0.000 0.011
DAB-DETR [28] ResNet50 28.939G 43.702M 0.002 0.002 0.000 0.005 0.346 0.769 0.219 0.762
Two-stage
Faster R-CNN [29] ResNet50 0.759T 33.035M 0.298 0.346 0.269 0.344 0.416 0.401 0.478 0.524
Cascade R-CNN [30] ResNet50 90.978G 69.152M 0.325 0.370 0.323 0.377 0.780 0.818 0.765 0.803
Grid R-CNN [31] ResNet50 0.177T 64.467M 0.299 0.402 0.306 0.393 0.746 0.826 0.720 0.786
Libra R-CNN [32] ResNet50 63.990G 41.611M 0.323 0.395 0.291 0.351 0.771 0.805 0.711 0.772
TridentNet [33] ResNet50 0.759T 33.035M 0.311 0.341 0.288 0.352 0.672 0.757 0.631 0.701
SABL [34] ResNet50 0.125T 42.213M 0.298 0.340 0.296 0.401 0.681 0.750 0.730 0.812
Dynamic R-CNN [35] ResNet50 63.179G 41.348M 0.314 0.355 0.309 0.372 0.761 0.805 0.659 0.738
One-stage
SSD [36] VGG16 87.552G 23.746M 0.296 0.427 0.279 0.480 0.847 0.919 0.842 0.947
RetinaNet [37] ResNet50 52.203G 36.330M 0.368 0.564 0.298 0.554 0.825 0.933 0.842 0.936
YOLOv3 [38] Darknet 50.002G 61.949M 0.200 0.311 0.158 0.342 0.779 0.918 0.830 0.881
CenterNet [39] ResNet50 50.278G 32.111M 0.201 0.497 0.202 0.544 0.853 0.924 0.847 0.916
FCOS [22] ResNet50 50.291G 32.113M 0.307 0.426 0.303 0.446 0.841 0.907 0.842 0.924
ATSS [40] ResNet50 51.504G 32.113M 0.355 0.429 0.355 0.475 0.849 0.911 0.861 0.907
CentripetalNet [41] HourglassNet 0.491T 0.206G 0.232 0.333 0.245 0.375 0.850 0.918 0.855 0.911
AutoAssign [42] ResNet50 50.555G 36.244M 0.326 0.416 0.336 0.433 0.842 0.918 0.851 0.925
GFL [43] ResNet50 52.296G 32.258M 0.269 0.337 0.356 0.478 0.857 0.917 0.864 0.924
PAA [44] ResNet50 51.504G 32.113M 0.356 0.668 0.342 0.573 0.957 0.941 0.868 0.933
VFNet [45] ResNet50 48.317G 32.709M 0.360 0.427 0.362 0.456 0.848 0.908 0.854 0.912
PVT-T [46] PVT 41.623G 21.325M 0.360 0.546 0.360 0.503 0.843 0.931 0.855 0.933
YOLOF [47] ResNet50 25.076G 42.339M 0.216 0.304 0.146 0.276 0.341 0.441 0.281 0.429
YOLOX [48] CSPDarknet 8.578G 8.968M 0.367 0.533 0.316 0.476 0.850 0.917 0.844 0.911
TOOD [7] ResNet50 50.456G 32.018M 0.297 0.390 0.286 0.427 0.860 0.914 0.864 0.908
DyHead [49] ResNet50 27.866G 38.890M 0.350 0.431 0.296 0.389 0.846 0.925 0.849 0.923
DDOD [50] ResNet50 46.514G 32.378M 0.281 0.362 0.338 0.456 0.855 0.921 0.856 0.928
RTMDet [51] CSPNeXt 51.278G 52.316M 0.276 0.476 0.216 0.450 0.332 0.765 0.353 0.760
EfficientDet [52] EfficientNet 34.686G 18.320M 0.279 0.442 0.359 0.560 0.849 0.934 0.829 0.911
$\star$ BAFE-Net (Ours) ResNet50 71.639G 35.626M 0.410 0.545 0.370 0.497 0.860 0.900 0.868 0.938
TABLE IV: Comparison with Other State-of-the-art infrared target detection methods on DenseSIRST, IRSTD-1k and IRSDSS.
Method DenseSIRST IRSTD-1k IRSDSS
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$ mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$ mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  ACM [11] 0.116 0.163 0.115 0.060 0.187 0.307 0.146 0.210 0.475 0.560 0.460 0.452
DNANet [13] 0.188 0.219 0.135 0.139 0.161 0.245 0.196 0.220 0.705 0.807 0.786 0.789
AGPCNet [14] 0.154 0.183 0.171 0.139 0.222 0.268 0.212 0.255 0.532 0.645 0.560 0.590
UIUNet [53] 0.178 0.235 0.208 0.198 0.349 0.475 0.192 0.175 0.771 0.761 0.831 0.826
MTUNet [54] 0.272 0.337 0.197 0.152 0.125 0.289 0.242 0.344 0.627 0.737 0.709 0.707
RDIAN [55] 0.101 0.124 0.172 0.128 0.107 0.149 0.115 0.138 0.650 0.754 0.694 0.731
URANet [56] 0.201 0.224 0.185 0.155 0.181 0.570 0.200 0.571 0.694 0.794 0.807 0.832
ABCNet [57] 0.261 0.315 0.189 0.159 0.335 0.406 0.202 0.196 0.684 0.842 0.763 0.786
SeRankDet [58] 0.216 0.261 0.114 0.120 0.308 0.408 0.160 0.268 0.658 0.742 0.774 0.788
$\star$ BAFE-Net (Ours) 0.274 0.342 0.248 0.338 0.410 0.545 0.370 0.497 0.860 0.900 0.868 0.938
V Experiments
V-A Experimental Settings
1) Dataset: During the experimental phase of our research, we conducted tests on the proposed dataset DenseSIRST. In addition to our proposed dataset, we also employed the publicly available IRSTD-1k dataset to further validate the effectiveness and generalization capability of our method.
2) Implementation Details: Considering the image sizes in the dataset and computational overhead, for methods within the DeepInfrared toolkit, we set the range of the long edge of training and testing images to [512, 512] while maintaining aspect ratio to avoid image distortion. We choose Focal Loss as the loss function of the classification branch, IoU Loss as the loss function of the regression branch, and Cross Entropy Loss as the loss function of the segmentation branch, with equal weighting assigned to three losses to ensure balanced multi-task optimization. We employ DAdaptAdam, a parameter-free adaptive optimization algorithm that dynamically adjusts learning rates based on gradient statistics, thereby improving training stability and reducing the need for manual learning rate tuning. The base learning rate is set to 1.0, and the weight decay is set to 0.05. To balance training time and convergence, we train all models for 24 epochs within the DeepInfrared framework.
@@ V-A Experimental Settings
TABLE V: Ablation study on the impact of Segmentation Head, Feature Exchange and Adapter on detection performance in Background-Aware Feature Exchange
@@
TABLE VI: Ablation study on the impact of Hidden Layers on detection performance in Adapter
@@
  (a) 8 0.243 0.327 0.206 0.308
(b) 16 0.252 0.337 0.214 0.310
(c) 32 0.254 0.326 0.218 0.312
(d) 64 0.260 0.329 0.230 0.323
(e) 128 0.254 0.326 0.211 0.333
(f) 256 0.243 0.319 0.212 0.312
TABLE VII: Ablation study on the impact of Exchange Method on detection performance in Feature Exchange Mechanism
@@
TABLE VIII: Ablation study on the impact of feature selection on detection performance in Channel-wise Feature Exchange
@@
TABLE IX: Ablation study on the impact of the number of exchange channels on detection performance in Top-K selection
@@
TABLE X: Ablation study on the impact of different ResNet on detection performance in BAFE-Net
Strategy ResNet DenseSIRST IRSTD-1k
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$ mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
  (a) ResNet18 0.283 0.335 0.233 0.325 0.433 0.603 0.380 0.528
(b) ResNet34 0.281 0.354 0.226 0.332 0.385 0.495 0.381 0.564
(c) ResNet50 0.274 0.342 0.248 0.338 0.410 0.545 0.370 0.497
(d) ResNet101 0.261 0.317 0.241 0.336 0.385 0.520 0.378 0.561
V-B Comparison with State-of-the-Arts
We benchmarked our BAFE-Net against current leading or state-of-the-art methods on the proposed DenseSIRST dataset, using metrics such as $mAP_{07}$, $mAP_{12}$, and Recall. The experimental results for general object detection frameworks are presented in Table II.
End-to-end (End2End) methods, while theoretically appealing due to their NMS-free style, demonstrated suboptimal performance on both datasets. This can be attributed to their inherent requirement for extensive training data, akin to COCO-level volumes, for effective operation. The relatively limited availability of infrared small target detection datasets inherently handicaps such methods, resulting in their diminished performance.
In contrast, one-stage methods, which directly regress bounding boxes and categories on the feature pyramids, demonstrated superior performance on both datasets. These methods are not only faster and more computationally efficient, but also achieved higher mAP and recall, indicating their overall effectiveness. Significantly, our BAFE-Net method outperformed all other one-stage methods on both datasets.
Beyond the DenseSIRST dataset, we further conducted experiments on the IRSTD-1k[5] and IRSDSS[59] datasets, as summarized in Table III. The results on IRSTD-1k were consistent with those observed on DenseSIRST, again highlighting the advantage of BAFE-Net in handling dense small-target detection. In contrast, the IRSDSS dataset mainly consists of maritime targets with relatively large object sizes. Consequently, while BAFE-Net still ranked among the best-performing approaches, the differences among the top methods on this dataset were relatively minor. Notably, BAFE-Net achieved state-of-the-art performance, with mAP scores matching or surpassing those of the best competing methods.
The superior performance of BAFE-Net consistently validates the efficacy of our approach in accurately detecting small infrared targets under diverse detection scenarios, including both densely clustered and sparsely distributed configurations.
Furthermore, to comprehensively situate BAFE-Net within the specific domain of infrared small target detection (IRSTD), we conducted comparisons against state-of-the-art methods specifically designed for this task. The results of these comparisons on both DenseSIRST, IRSTD-1k and IRSDSS datasets are presented in Table IV. As shown, BAFE-Net also achieves superior performance compared to these specialized IRSTD methods, further demonstrating its effectiveness.
@@ V-C Visual Analysis
Figure 5: Feature visualization comparison. This comparison showcases the detection results of different methods in terms of feature visualization. The proposed BAFE-Net demonstrates superior performance by effectively detecting small objects with fewer false alarms.
V-D Ablation Study
V-D1 Ablation Study on the Scheme of Background-Aware Feature Exchange
The BAFE module constitutes the core contribution of this paper, as evidenced by its significant impact on detection performance demonstrated in Table V. To validate the effectiveness of the proposed components, we conducted an extensive ablation study by exploring various combinations of the background-based segmentation, Feature Exchange module, and adapter module.
Table V summarizes the results and reveals several key insights. First, each individual module contributes significantly and independently to detection performance, thereby validating the foundational design of our framework. Second, a comparison between Strategy (a) and Strategy (b) highlights the benefit of incorporating background-based segmentation into the object detection pipeline, where the detection head interacts with the segmentation head. This integration results in performance gains of 2.1% and 0.3% on the DenseSIRST dataset.
Furthermore, the performance difference between Strategy (b) and Strategy (c) demonstrates the effectiveness of the Feature Exchange mechanism in enhancing the interaction between the detection and segmentation heads. This enhancement leads to improvements of 0.8% and 1.6% on the DenseSIRST dataset. Finally, comparing Strategy (c) with Strategy (d) shows that introducing a feature adapter prior to the Feature Exchange further amplifies detection performance.
As the feature adapter employs a Multi-Layer Perceptron (MLP), the number of hidden layers in the MLP is a crucial parameter that warrants careful consideration. Increasing the number of hidden layers typically improves performance but comes at the cost of increased computational complexity. To strike a balance between performance and efficiency, we conducted an experiment to determine the optimal number of hidden layers, as shown in Table VI. The results show that performance consistently improves as the number of hidden layers increases, up to 64 layers. However, when the number of hidden layers exceeds 64, the detection performance begins to decline, while the parameter count continues to grow significantly. Therefore, we select 64 hidden layers for the Adapter, as it offers the best trade-off between detection accuracy and computational overhead.
V-D2 Ablation Study on Feature Exchange Mechanism
In this section, we conduct a rigorous investigation of various Feature Exchange Mechanisms, including Channel Exchange, Spatial Exchange, and several of their combinations. The goal is to identify the most effective strategy for facilitating cross-task feature interaction. Channel Exchange enables direct information flow between features at the channel level, while Spatial Exchange operates by exchanging information along the spatial dimensions of the feature maps. To ensure an equitable exchange of features, the exchange ratio is consistently fixed at 0.5 in all configurations.
As shown in Table VII, Channel Exchange consistently outperforms other Feature Exchange strategies. The experimental results clearly highlight its effectiveness and stability, confirming its selection as the optimal mechanism for feature exchange within our model.
V-D3 Ablation Study in Dynamic Channel Selection
In this section, we perform an in-depth investigation of different feature selection strategies, including random selection, fixed selection, and dynamic selection, with the goal of identifying the most effective approach for enhancing feature exchange. 1) Randomly Selection: This method selects feature channels randomly, which may lead to suboptimal feature exchanges. In some cases, it could even select features that hinder the task, ultimately reducing model performance. 2) Fixed Selection: While fixed selection provides consistency during training and eliminates uncertainty, it lacks the adaptability to dynamic input features, thus limiting the model’s performance in complex or varied scenarios. 3) Dynamic Selection: dynamic selection maximizes the performance of the model and reduces unnecessary information redundancy by selecting the most useful feature channels for exchange. The exchange ratio is judiciously set to 0.5 to maintain a balance.
As shown in Table VIII, the advantages of dynamic selection are obvious. Empirical evidence highlights the robustness and optimality of dynamic selection, thus confirming its adoption as the preferred method for feature selection in our model.
Additionally, we conduct a focused analysis on the impact of the hyperparameter $p$ in the Top-K selection mechanism. As shown in Table IX, setting $p=1/2$ yields the best performance, validating this choice as optimal for our design.
V-D4 Impact of Backbone Depth
In infrared small target detection, the depth of the backbone network plays a pivotal role in determining detection performance. A shallower backbone, though computationally efficient, may fail to capture sufficiently rich semantic features, thereby limiting the model’s ability to distinguish small targets from complex backgrounds. On the other hand, an overly deep backbone may introduce excessive abstraction, potentially suppressing the subtle features of small targets and leading to performance degradation. Therefore, achieving an optimal balance between depth and feature representation is essential for effective infrared target detection.
We conducted a series of experiments using different ResNet architectures, varying the depth from ResNet18 to ResNet101, to identify the optimal backbone for our proposed BAFE-Net. The results of these experiments are summarized in Table X. Notably, on DenseSIRST, ResNet18 has the highest $mAP_{07}$ metric and ResNet50 has the highest $mAP_{12}$ metric; on IRSTD-1k, ResNet18 has the highest $mAP_{07}$ metric and ResNet34 has the highest $mAP_{12}$ metric.
@@ V-D4 Impact of Backbone Depth
Second, our method generates cleaner background activation maps. As evident from Fig. 5, the background regions in our activation maps exhibit consistently lower activation intensities, effectively suppressing background interference. This can be attributed to the background-aware mechanism employed by BAFE-Net, which learns to differentiate between target and background regions during the feature exchange process. By minimizing the activation levels in the background areas, our method demonstrates enhanced robustness against false positives, leading to a higher detection precision.